The sum of the ages of David and Martha is 60 years, the difference is 16 years. If David is older than Martha, which of these statements are correct?
A) David is 38
B) Martha is 38
C) David is 44
D) Martha is 44
I know the answer is A. But how do I solve the problem to get the answer.

Answer by stanbon(75887)

The sum of the ages of David and Martha is 60 years, the difference is 16 years. If David is older than Martha, which of these statements are correct?
---------
Equations:
D + M = 60
D - M = 16
-----
Add and solve for "d":
2D = 76
D = 38 yrs (David's age)
-----
Solve for M::
D + M = 60
38 + M = 60
M = 22 yrs (Martha's age)
====================================
Cheers,
Stan H.
------------------------

A) David is 38
B) Martha is 38
C) David is 44
D) Martha is 44